3. Consider the following linear programming problem: Maximize 10X + 12Y Subject to: 8X + 4Y ≤ 840 2X + 4Y ≤ 240 X, Y ≥ 0 Graph the constraints and shade the area that represents the feasible region. Find the solution to the problem using either the corner point method or the isoprofit method. What is the maximum feasible value of the objective function?
Answer #1

la (0,2107 Q. 8x + 4y < 840 :. 2x+Y< 210-- 2x+4y=240 X+27 $120--0 X,Y>0---- Z = 10x+12Y 1 (0,60, 2x+=210 100,10) - X+2Y=120 V